OpenStudy (anonymous):

an equation of the line containing the given point and parallel to the given line (4.5)x+2y=9

OpenStudy (angela210793):

2y=9-x y=9/2-x/2 slope=-1/2 eq of the line ur asking is: y-5=-1/2(x-4) y-5+x/2-2=0 x/2+y-7=0

OpenStudy (anonymous):

im lost what would be the equation of the line y=???

OpenStudy (angela210793):

x/2+y-7=0 or y=-x/2+7

OpenStudy (amistre64):

what is the given point?

OpenStudy (angela210793):

(4;5) i guess